OpenStudy (mkmkasim):

does ∫from 1 to 2 of (1/(x-1^4)) converge or diverge?

OpenStudy (anonymous):

\[\int_1^2\frac{1}{(x-1)^4}~dx~~?\]

OpenStudy (anonymous):

I only ask because you could be asking \[\int_1^2\frac{1}{x-1^4}~dx=\int_1^2\frac{1}{x-1}~dx\]